equivalent expression with out parentases -(-9+8y)
Answer by drj(1380)

Equivalent expression with out parenthesis -(-9+8y)

Step 1. Think of the - sign outside the parenthesis as -1 and then distribute.

Step 2. -(-9+9y)=(-1)*(-9+8y)=(-1)*(-9)+(-1)*(8y)=9-8y

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J